the measure of an angle is 15 less than twice the measure of its complement. Find the measure of each angle
Answer by cleomenius(959)

x - 15 = 2(90 - x)
x - 15 = 180 - 2x
3x = 165
x = 55, the angle, the complement is 35.
35 * 2 = 70; 70 - 15 = 55.
This does check.
Cleomenius.